# Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training

### v1

@@ Abstract
abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications. To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy. However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment. To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity. Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
Vision-Language Pre-training, Federated Learning, Data Heterogeneity
abstract Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for biomedical applications. To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for biomedical VLP while protecting data privacy. However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment. To address this challenge, we propose Federated distributional Robust Guidance-Based (FedRGB) learning framework for federated VLP with robustness to data heterogeneity. Specifically, we utilize a guidance-based local training scheme to reduce feature distortions, and employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. The experiments on real-world datasets show our method successfully promotes efficient federated multimodal learning for biomedical VLP with data heterogeneity.
Figure 1: Federated multi-modal pre-training aims to de-centralizedly utilizing private dataset to tackle the data hungry problem. Fitting heterogeneous local datasets that correspond to different distributions of latent semantics (e.g. disease) would result in biased models. Our FedRGB framework has successfully upgraded multi-modal pretraining methods by introducing robust alignment module as guidance during local training time.
@@ 1 Introduction
Vision-language pre-training (VLP) learns efficient multimodal representations by extracting latent semantics from a large amount of image-text pairs, where the dataset scale largely impacts the performance of the learned model (Oquab et al., 2023). However, enlarging pre-training dataset is a non-trivial challenge especially for biomedical applications which is restricted by privacy issues of clinical and patient data (Ladbury et al., 2023). To utilize data across different institutes, recent work (Lu et al., 2023a) has attempted to introduce federated learning into the biomedical VLP domain. However, in real-world scenarios, datasets collected from different institutes are always heterogeneous, which is a long-standing practical challenge in federated learning (Ghosh et al., 2019; Zhang et al., 2023a; Huang et al., 2022). For example, one hospital receives a high proportion of patients with pneumonia, while another hospital may undertake more patients with tuberculosis, due to variety of patients cohort and geographics. While the data heterogeneity would essentially damage the performance of the federally trained model, current FL methods for biomedical VLP (Lu et al., 2023a) haven’t taken this vital factor into consideration to provide an effective solution. To address this issue, simply averaging client models trained on heterogeneous local datasets results in biases as shown in Figure 1. Firstly, the client model would be heterogeneous, and re-training on the sent-back server model might distort the aggregated structure of the representation space. Secondly, averaging model parameters cannot produce a robust alignment between the two modalities, which not only affects the generalization ability, but also prevents the model from learning more diverse correlations between vision and language. Furthermore, as we will demonstrate in our empirical findings (Sec 5.2), heterogeneity harms the performance of pre-trained model by affecting the deep layers which serve as an alignment module in the VLP model. In this paper, we propose the FedRGB framework to mitigate heterogeneity in federated biomedical VLP to enable its deployment in real-world scenarios. To alleviate the distortion in local client training, we introduce a teacher alignment module to provide unbiased cross-modal alignment by generating anchor representation space for regularization. We employ a distributionally robust optimization (DRO)-based algorithm to learn a robust teacher alignment module, which is optimized to better alleviate the worst-case distortion on the feature encoder caused by local cross-modal alignment. Our contributions primarily focus on:
@@
  • We for the first time tackle the problem of biomedical VLP under the federated setting by utilizing heterogeneous datasets from different institutes. We conduct empirical studies to analyze the influence of data heterogeneity on federated multimodal learning.
@@
  • We propose FedRGB framework to mitigate the heterogeneity in federated multimodal learning. Specifically, our method introduces the guidance-based local training scheme for learning unbiased cross-modal alignment to alleviate representaion distortions.
@@
  • Experiment results show the effectiveness of our method in learning multimodal representations under the federated setting by appling to various downstream tasks including image-text retrieval, classification, and segmentation.
@@ 2.1 Vision-Language Pre-training
Vision-Language Pre-training (VLP): Leveraging data of different modalities to enable multimodal reasoning and representation learning for various cross-modal tasks has become an active research field in machine learning (Radford et al., 2021). Pre-training vision-language models on large datasets and then transferring learned knowledge to downstream tasks has become a popular approach to extract semantics from unlabeled image-text paired datasets (Li et al., 2022a; Bao et al., 2022). To find the shared latent space, current works often focus on learning generalizable and robust alignment relationships between the two modalities (Radford et al., 2021; Chen et al., 2020a), which measures the ability to connect and retrieve the representation of the paired data. Biomedical Vision-Language Pre-training: Most of the existing works in biomedical vision-language pre-training are based on the twin modality-specific encoders for medical images and reports respectively, with the image-report contrastive loss (Zhang et al., 2022b) to align visual and language representations on a joint latent space. A significant improvement has been achieved using the local contrastive loss which aims to align fine-grained representations (Huang et al., 2021). Incorporating medical domain knowledge, Wang et al. (2022) proposes a well-designed method that aligns the image and report embeddings in three different granularities. Recent works have also tried to address other practical issues of misalignments (Bannur et al., 2023). However, most of these methods haven’t considered the adaptation to de-centralized scenarios during pre-training and thus might fail to handle the potential damage from data heterogeneity.
2.2 Federated Learning
Heterogeneous Federated Learning: Federated learning aims to de-centralize model training by aggregating models trained on local clients to obtain more generalizable models with scaled-up training datasets while protecting data privacy (McMahan et al., 2017), which achieved success in many applications especially those with data privacy issues like biomedical data. However, client datasets are heterogeneous in most cases, which damages the performance of the aggregated model (Ghosh et al., 2019). Existing methods have tried to solve the data heterogeneity problem in many ways (Li & Wang, 2019; Collins et al., 2021). However, most methods either require additional communications on representations (Zhang et al., 2023a) which are not privacy-preserved and inefficient, or need large-scale public datasets for server aggregation (Huang et al., 2022; Yu et al., 2023b) that does not fit the case for biomedical data. Besides, remaining methods are often designed for supervised tasks that focus on label distribution (Diao et al., 2020) or feature distribution (Li et al., 2022b), which cannot be directly adapted to the vision-language pre-training task. While the self-supervised pre-trained models can be biased due to the heterogeneous datasets in the biomedical domain (Krishnan et al., 2022) and over-reliance on spurious features (Saab et al., 2022), distributionally robust optimization (Deng et al., 2020) can optimize the worst-case performance thus alleviating these issues (Liu et al., 2022). Recent works have considered its application in unsupervised and self-supervised learning senarios (Han et al., 2023; Zhang et al., 2022a; Capitani et al., 2024). Federated Multi-modal Learning: Federated learning strategies have been extended to multimodal datasets in recent works (Zhou & Wang, 2022), with the goal to reduce training cost and promote personalized deployment of multimodal foundation models (Su et al., 2022; Chen et al., 2023b; Lu et al., 2023b). For example, Yu et al. (2023a) considers a client-server collaborative pre-training setting, benefiting the pre-training process of the server model by leveraging representations from local modalities. Nevertheless, these works require a pre-training stage in the server while ignoring that only on client datasets with limited samples. The closest work to ours is Lu et al. (2023a), which adopts federated learning to scale up the pre-training dataset. However, this work lacks generalization ability and hasn’t considered the data heterogeneity issue that is very common and significant in practice.
@@ 3 Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training biomedical vision-language models to better capture the semantic alignment between the two modalities, with heterogeneous private datasets from different clients. In this section, we will formulate our problem set-up and assumptions. Please see Appendix A for the detailed preliminary. Federated Vision-Language Pre-training. In this paper, we consider the federated vision-language pre-training (VLP) problem in the biomedical domain. Suppose we are given $N$ clients $\{C_{i}\}_{i=1}^{N}$, each client has a corresponding local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy issues. Without access to centralized pre-training, federated pre-training aims to conduct vision-language representation learning across local clients, while still capture a robust and generalizable cross-modal alignment, by efficiently aggregating client models $\{M_{i}\}_{i=1}^{N}$ trained on their private datasets. Similar to conventional federated learning scenarios, our federated pre-training consists of multiple communication turns. For communication turn $t$, client models $\{M_{i}\}_{i=1}^{N}$ completed the same number of local iteration steps respectively, and the server receives and aggregates the weights sent from local models to produce the aggregated server model $M^{t}$. Then it will be sent back to clients and re-trained on $\{D_{i}\}_{i=1}^{N}$ separately to get new local models $\{M_{i}^{t+1}\}_{i=1}^{N}$. The federated vision-language pre-training is a more practical and general task compared to related works. Table 1 lists the most similar settings in previous works, and the difference between our task and theirs.
Table 1: Comparison between related multi-modal federated learning tasks and our federated biomedical vision-language pre-training task. We organize the related works according to our key areas of interest. (1). Heterogeneous clients. (2). Multi-modal clients. (3). Transferability. (4). Prior model-free. (5). No communication on public data or feature representations.
Task hetero. m.m.c. transf. no prior no feat. com.
(Yu et al., 2023a) $\checkmark$ $\checkmark$
(Lu et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$
(Chen et al., 2023b) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
(Lu et al., 2023a) $\checkmark$ $\checkmark$ $\checkmark$
FedRGB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Data Formulation. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of respective data samples $X=\{x_{i}\}_{i=1}^{n},Y=\{y_{i}\}_{i=1}^{m},n,m\in N^{+}$. We assume there exists latent variables $U$ that embed semantics across two modalities. For example, the disease category of a given X-ray radiology-report pair is a latent variable connecting these two modalities, which determines the pathology region of the radiology image and semantics in the diagnosis report. For heterogeneous local datasets, we assume the distribution of latent variables $P_{i}(U)$ is different across clients.
@@ 4 Method
Figure 2: (a) Our proposed FedRGB upgrades backbone multi-modal pre-training strategies via introducing a robust guidance to local pre-training. This approach can reduce the distortion on feature encoders and learn better cross-modal alignment. (b) We employ a distribution-based min-max optimization to learn robust cross-modal alignment, which will be used as guidance in local pre-training.
We propose a de-centralized framework to mitigate the data heterogeneity in federated biomedical vision-language pre-training, for learning an efficient vision-language model that can robustly capture reliable cross-modal semantic alignments. In this section, we first introduce our approach with a guidance-based local training strategy to reduce the distortion on encoders caused by fitting client datasets. Secondly, we will illustrate how our approach employ a distribution-based min-max optimization to learn unbiased cross-modal alignment. Algorithm 1 shows the pipeline of our method.
@@
4.1 Guidance-based Local Client Training
As shown in our empirical findings (Sec. 5.2), biomedical vision-language models pre-trained on heterogeneous client datasets are affected by the local correlations of the two modalities, which can’t be addressed by simply averaging the weights of local models. Furthermore, this mainly falls on the deep layers which act as alignment modules, where the heterogeneous alignment module would distort the learned encoder during local iterations as shown in Sec. 7. Inspired by these findings, we introduce a teacher aligner to guide the representation space learned by local models to converge to a generalizable and robust one. Teacher alignment module. To alleviate the distortion and bias when learning local features encoders and alignment modules, we propose to utilize a teacher alignment module to generate anchor representations conditioned on the output feature vector of the encoder to encourage local models to learn a homogeneous representation space across clients. Suppose after one client-server communication turn, given the local dataset of client $c\in\mathcal{C}$ sampled from $(x^{c},y^{c})\sim D_{c}$, and teacher alignment module $g_{v}^{*},g_{l}^{*}$ corresponding to vision and language modality, we target at updating the feature encoders of image and text $f_{v},f_{l}$ and the corresponding alignment modules $g_{img},g_{txt}$, by training on the local client dataset $D_{c}$, under the guidance of teacher alignment modules. Guiding local training with robust aligner. As shown in Fig. 2, given the image-text pairs from local dataset as input $(x^{c},y^{c})\sim D_{c}$, we denote the feature vectors output from the feature encoder modules as $(v^{c},l^{c})$. Then getting through the alignment module, the aligned representations output from local aligners are denoted by $(z_{I}^{c},z_{T}^{c})$, while those from teacher aligners are $(z^{c*}_{I},z^{c*}_{T})$. The goal is to reduce the discrepancy between two sets of aligned representations produced by local and teacher aligners, so that we can keep locally trained model will not be misled by the bias from local datasets and all local clients can converge to learn a similar embedding space that captures globally reliable and generalizable cross-domain alignment. The teacher guidance loss can be formulated as below:
$\footnotesize\mathcal{L}_{guide}=\sum_{i=1}^{N_{c}}(z^{c*}_{I,i}-z^{c}_{I,i})^{2}+\sum_{i=1}^{N_{c}}(z^{c*}_{T,i}-z^{c}_{T,i})^{2}$ (1)

### v2

# Align as Ideal: Cross-Modal Alignment Binding for Federated Medical Vision-Language Pre-training

@@ Abstract
Vision-language pre-training (VLP) has arised as an efficient scheme for multimodal representation learning, but it requires large-scale multimodal data for pre-training, making it an obstacle especially for medical applications. To overcome the data limitation, federated learning (FL) can be a promising strategy to scale up the dataset for medical VLP while protecting data privacy. However, client data are often heterogeneous in real-world scenarios, and we observe that local training on heterogeneous client data would distort the multimodal representation learning and lead to biased cross-modal alignment. To address this challenge, we propose a Federated Align as IDeal (FedAID) framework for federated VLP with robustness to data heterogeneity, to bind local clients with an ideal cross-modal alignment. Specifically, to reduce distortions on global-aggregated features while learning diverse semantics from client datasets during local training, we propose to bind the cross-model aligned representation space learned by local models with an unbiased one via guidance-based regularization. Moreover, we employ a distribution-based min-max optimization to learn the unbiased cross-modal alignment at each communication turn of federated pre-training. The experiments on real-world datasets demonstrate our method successfully promotes efficient federated multimodal learning for medical VLP with data heterogeneity.
@@ 1 Introduction
Vision-language pre-training (VLP) learns efficient multimodal representations by extracting latent semantics from a large amount of image-text pairs, where the dataset scale largely impacts the performance of the learned model (Oquab et al., 2023). However, enlarging multimodal pre-training dataset is a non-trivial challenge especially for medical applications which is restricted by privacy issues of clinical and patient data Ladbury et al. (2023). To collect and utilize data across different institutes, recent work Lu et al. (2023a) has attempted to introduce federated learning into the medical VLP domain.
However, in real-world scenarios, datasets collected from different institutes are always heterogeneous, which is a long-standing practical challenge in federated learning Ghosh et al. (2019); Zhang et al. (2023a); Huang et al. (2022). For example, one hospital receives a high proportion of patients with pneumonia, while another hospital may undertake more patients with Edema, due to the variety of patient cohorts and geographics. While the data heterogeneity would essentially damage the performance of the federally trained model, current medical multi-modal pre-training methods haven’t taken this vital factor into consideration, which encounters the deployment of medical VLP into decentralized pre-training scenarios. For a thorough study especially under the case of multi-modal learning, while multi-modal pre-training aims to learn a well-aligned latent space shared by each modality Zhang et al. (2022b); Li et al. (2021a), this cross-modal alignment would be affected by the data heterogeneity issue during local training Zhang et al. (2021); Ye et al. (2023), thus failing to provide an effective solution.
The current solution of simply averaging client models trained on heterogeneous local datasets is sub-optimal as shown in Fig. 1. First of all, when the client datasets are heterogeneous, re-training on the sent-back server model might distort the aggregated structure of the representation space. Secondly, as observed in empirical study Sec 5.2, averaging model parameters cannot produce an ideal alignment between two modalities which captures robust and factual diverse correlations between vision and language. And biased cross-modal alignment which overfits multi-modal correlations of local datasets, will not only affect generalization ability, but prevent the model from learning transferable and diverse semantics from local datasets. Furthermore, as we demonstrate in our empirical findings (Sec 5.2), heterogeneity harms the performance of pre-trained models by affecting the deep layers, which serve as alignment modules in the VLP model.
Figure 1: Federated multi-modal pre-training aims to decentralizedly utilize private datasets to tackle the data-hungry problem in pre-training transferable models. During local training, directly fitting heterogeneous local datasets that w.r.t. different distributions of latent semantics (e.g. disease) would result in non-ideal cross-modal alignment. Our method upgrades multi-modal pretraining methods to federated setting by binding local cross-modal alignment to an ideal one during local pre-training.
In this paper, we propose Federated Align as IDeal (FedAID) framework to mitigate data heterogeneity effect on encountering learning ideal cross-modal alignment in federated vision-language pre-training and enable its deployment in real-world scenarios. To alleviate the distortion on global-aggregated features during local client training while still encouraging models to extract diverse semantics from local datasets, we bind the cross-modal aligned representation space learned by local models to an unbiased representation space. Specifically, we introduce a teacher alignment module to provide robust cross-modal alignment by generating anchor representation space for regularization. We employ a distributionally robust optimization (DRO)-based algorithm to learn a robust teacher alignment module, which is optimized to better alleviate the worst-case distortion on the feature encoder caused by locally heterogeneous cross-modal alignment.
Our contributions primarily focus on:
@@
  • We for the first time tackle the problem of medical VLP under the federated setting by utilizing heterogeneous multi-modal datasets from different institutes. We conduct empirical studies to analyze the influence of data heterogeneity on federated multi-modal learning.
@@
  • We propose FedAID to mitigate the data heterogeneity on encountering learning ideal cross-modal alignment in federated VLP. Specifically, our method binds representations aligned by client models with unbiased representation space by introducing the guidance-based local training for learning unbiased cross-modal alignment to alleviate feature distortions.
@@
  • Experiment results show the effectiveness of our method in learning multi-modal representations under the federated setting by appling to various downstream tasks including image-text retrieval, classification, and segmentation.
@@ 2 Related Work
Medical Vision-Language Pre-training. Pre-training vision-language models on large-scale datasets and then transferring learned knowledge to downstream tasks has become a popular approach to leverage diverse semantics contained in multi-modal unlabeled data  Li et al. (2022a); Bao et al. (2022); Radford et al. (2021). To obtain more generalizable representations, current works aim to learn the latent space shared by each modality using contrastive learning via cross-modal alignment  Radford et al. (2021); Chen et al. (2020) or multi-modal fusion Li et al. (2021a); Chen et al. (2022), which connect the representations from the paired data. Existing works in medical vision-language model pre-training employ twin modality-specific encoders for mapping medical images and reports to the latent space respectively, and use the image-text contrastive loss Zhang et al. (2022b) to align visual and language representations. Improvement has been achieved by using the word-patch contrastive loss which aims to align fine-grained cross-modal representations Huang et al. (2021a). Furthermore, incorporating medical domain knowledge,  Wang et al. (2022) proposes to align the image and report embeddings in three different granularities. Recent works have also tried to address other practical issues of misalignments Bannur et al. (2023). However, almost all of the current methods still rely on large-scale pre-training datasets, which impede the adoption to modalities with limited training samples and deployment in real-world scenarios.
Federated multi-modal learning. Federated multi-modal learning Zhou & Wang (2022) aims to facilitate diverse semantics in local datasets in a decentralized and privacy-preserved wayRelated works often focus on leveraging client-side information to train personalized information while utilizing server-side pre-trained models.  Su et al. (2022); Chen et al. (2023); Lu et al. (2023b), and use local datasets to enhance the centralized training Yu et al. (2023). However, these works haven’t answered a key question: how can we decentralize the multi-modal pre-training through a federated way in real-world scenario to learn generalizable cross-modal representations? Recent advances Lu et al. (2023a) has verified that federated learning can be utilized to scale up the pre-training dataset. However, this work hasn’t considered the harm of data heterogeneity issue Ghosh et al. (2019), which is very common and significant in practice such as in medical applications.
Heterogeneity in federated learning. Existing methods have tried to solve this problem in uni-modal supervised learning tasks Li & Wang (2019); Collins et al. (2021); Li et al. (2021b). These methods designed for supervised tasks focus on the heterogeneous problem in label distribution Diao et al. (2020) or feature distribution Li et al. (2022c), which cannot be directly adapted to the vision-language pre-training task. Recent works in representation-augmentation based uni-modal self-supervised federated learning Zhang et al. (2023a); Huang et al. (2022); Li et al. (2022b) employ additional communications on representations of local data to increase sample diversity. Such methods fail to protect data privacy, which is a vital concern in medical applications. Yan et al. (2023) utilizes image-side-only techniques to handle heterogeneity and thus may harm the learning of cross-modal alignment when adapted to the multi-modal setting.  Zhuang et al. (2022); Li et al. (2021b) considers distillation-based methods yet ignores the direct modeling of cross-modal alignment. While the learned models can be biased and over-reliance on spurious correlations Saab et al. (2022) due to the data heterogeneity in the medical domain Krishnan et al. (2022), distributionally robust optimization Deng et al. (2020) can alleviate these issues by optimizing the group-wise worst-case performance on given objective Liu et al. (2022), thus can be flexibly adapted to various of federated learning tasks Han et al. (2023); Zhang et al. (2022a); Capitani et al. (2024).
@@ 3 Problem Formulation
We aim at mitigating the data heterogeneity issues in federally pre-training medical vision-language models from heterogeneous private client datasets. In this section, we will formulate the set-up of the Federated Vision-Language Pre-training task, which aims to capture generalizable cross-modal alignment by leveraging diverse semantics from local datasets. Then we will introduce our assumption on multi-modal pre-training data and heterogeneity. Please see Appendix F for detailed preliminary.
Federated Vision-Language Pre-training (FedVLP). We aim to federally pre-train medical vision-language models. Specifically, given $N$ clients $\{C_{i}\}_{i=1}^{N}$, FedVLP decentralizedly conducts vision-language representation learning across clients, where local models are learned only on the local client dataset. The goal is to aggregate the locally trained models on a central server to obtain a generalizable and unbiased global pre-trained model. Here, we assume each client has a local yet private dataset $\{D_{i}\}_{i=1}^{N}$ that can’t be easily shared due to privacy or other issues. During FedVLP, different from supervised federated learning, we focus on the transferability of the aggregated model. Different from supervised federated learning which is usually designed for a specific task, federated vision-language pre-training focuses on learning a generalizable and unbiased representation space, which can be transferred to various downstream tasks as mentioned in Sec 5.1.
Employing an alternative updating approach, federated pre-training also consists of multiple communication turns between clients and the server , which is the same as conventional federated learning (FL) settings. For each communication turn $t$, client models $\{M_{i}^{t}\}_{i=1}^{N}$ completed local pre-training respectively and send local model weights to server, where all the local model weights are aggregated to obtain a server model $M^{t}$. $M^{t}$ will be broadcast to clients for updating local models in each communication turn, which completes the main part of the communication turn. In some FL methods, additional local computation and information transmissions may be conducted after clients have received the aggregated model, which will also be counted as a part of the communication. Table 18 in appendix lists the most similar settings, and their corresponding differences with ours.
@@
Formulation of Multi-Model Pre-training Data and Heterogeneity. We consider the multimodal datasets including vision and language modalities $X,Y$, consisting of image and text data samples respectively, $X=\{x_{i}\}_{i=1}^{n_{1}},Y=\{y_{i}\}_{i=1}^{n_{2}},n_{1},n_{2}\in N^{+}$. We assume there exist $m$-way discrete multi-label latent variables $U\in R^{m}$ that embed semantics shared across two modalities, and form multiple latent classes $\{u_{i}\}_{i=1}^{k},k\in N^{+}$. For example, disease labels of a given X-ray radiology-report pair, are latent semantic variables that connect these two modalities, which determines the pathology region of the radiology image and corresponding description in the diagnosis report. To model heterogeneity, we assume the distribution of latent variables is different across clients.
Cross-modal alignment. We model cross-modal alignment as a module that outputs embeddings with high semantic similarity inspired by Arora et al. (2019). Therefore, we formulate the goal of the cross-modal alignment function, which is to minimize the distance between the distributions induced by vision and language representations. This could be achieved by employing a contrastive-based self-supervised learning object Huang et al. (2021b); Arora et al. (2019). Detailed analysis of this modeling can be seen in Appendix G.1.
@@ 4 Method
Figure 2: (1) Federated pre-training strategies with uni-modality in previous works Zhuang et al. (2022); Jiang et al. (2023) often consider the representation or performance discrepancy, adapting them directly to multi-modal pre-training problem underestimates the necessity of modeling multi-modal interactions. (2) Our method considers the cross-modal alignment and binds local alignment to an unbiased one during local training. (3) In each communication turn, after local training and aggregating encoders, We employ a distribution-based min-max optimization to learn robust cross-modal alignment (3.b, 3.c). It will be used to guide local pre-training (c.1).
To federally leverage heterogeneous local datasets to pre-train an efficient vision-language model, we propose a decentralized framework that can robustly capture reliable cross-modal semantic alignments across different clients. As shown in Fig. 2, we employ the teacher alignment module to guide the local training, to learn factual diverse correlations between two modalities.
In this section, we first introduce our approach to reduce distortion on global-aggregated features during fitting local dataset via binding cross-modal alignments, through a guidance-based local training strategy. Secondly, we will illustrate our approach of employing the distributionally robust optimization to learn unbiased cross-modal alignment. Details can be seen in Algorithm 1.
@@
4.1 Cross-Modal Alignment Binding via Guidance-Based Local Training
As shown in our empirical findings (Sec. 5.2), medical vision-language models pre-trained on heterogeneous client datasets are affected by the local correlations of the two modalities, which can’t be addressed by simply averaging the weights of local models. Furthermore, this mainly falls on the deep layers which act as the cross-modal alignment module. As shown in Sec. 8, we observe that a non-ideal alignment module would amplify bias in local datasets and enforce the encoder to learn biased features during local iterations and distort the global-aggregated features. Inspired by these findings, we design to separately handle the deep layers and shallow layers of the modality-specific network, which are considered to be the encoder and the alignment module, respectively. To encourage models to capture semantics in local datasets during local training, we keep updating the whole model during local training. To avoid feature distortion, we focus on the alignment module and bind the local cross-modal alignment to an ideal one.
Binding cross-modal alignments on the representation space. Binding local cross-modal alignments can be straightforwardly solved by replacing the local alignment module with a frozen unbiased one. However, this is not practical due to our goal of encouraging models to learn semantics from local datasets. Besides, the unbiased alignment module of the current communication turn should also be updated. To enable cross-modal alignment binding with fully learnable local models, we employ an alternative approach to minimize the distance between the representation space projected by local alignment modules and an anchor representation space.
As shown in Fig. 2, given input image-text pairs from dataset of client $c\in\mathcal{C}$ as $(x^{c},y^{c})\sim D_{c}$, where c is the index of client, the feature vectors output from the image and text encoder modules can be denoted as $(I^{c},T^{c})$, respectively. $(I^{c},T^{c})$ will be aligned by alignment modules, and the aligned representations output from local aligners and anchor representations output from unbiased aligner are denoted by $(z_{I}^{c},z_{T}^{c})$, $(z^{c*}_{I},z^{c*}_{T})$ respectively. During the local training, given the learning object of aligning two modalities $L_{task}(x,y)$ (e.g. image-text contrastive loss in ConVIRT Zhang et al. (2022b)), we have:
Remark 1
Consider $L$-lipschitz learning object $L_{\text{task}}:\mathbb{R}^{2d}\to\mathbb{R},L_{\text{task}}>0,d\in\mathbb{N}^{+}$. Minimizing $L_{\text{task}}(z^{c*}_{I},z^{c*}_{T})$ can be reformulated as: $\min_{\theta}\left(L_{\text{task}}(z_{I}^{c},z_{T}^{c})+L\left(\|z_{I}^{c}-z^{c*}_{I}\|_{2}^{2}+\|z_{T}^{c}-z^{c*}_{T}\|_{2}^{2}\right)\right)$, where $\theta$ represents the model parameters. See detailed analysis in Appendix G.1.
Guiding local training with robust alignment module. To bind local cross-modal alignments, motivated by Remark 1, we can introduce a robust teacher alignment module to generate anchor representations conditioned on the output feature vector of the encoder, and leverage them for guiding local training. Specifically, given the local dataset of client $c\in\mathcal{C}$ sampled from $(x^{c},y^{c})\sim D_{c}$, and teacher alignment module $g_{I}^{*},g_{T}^{*}$, we target at updating the feature encoders $f_{I},f_{T}$ and the corresponding alignment modules $g_{I},g_{T}$, by training on the local client dataset $D_{c}$, under the guidance of teacher alignment modules. Following the above analysis, we aim to reduce the discrepancy between two sets of aligned representations produced by local and teacher alignment modules. Thus, given number $N_{c}$ of samples in client $c$, the teacher guidance loss can be formulated as below:
$\scriptsize\mathcal{L}_{guide}=\sum_{i=1}^{N_{c}}(z^{*}_{I,i}-z_{I,i})^{2}+\sum_{i=1}^{N_{c}}(z^{*}_{T,i}-z_{T,i})^{2}.$ (1)